OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (anonymous):

The formula to calculate a specific term in a geometric sequence is: an = a1(r)^(n - 1) But, you don't know what the ratio is. So solve for the ratio using the formula. 25 = 625(r)^(3 - 1) 25 = 625(r)^(2) 25 / 625 = (r)^2 0.04 = r^2 0.2 = r So now you know what the ratio is, use the formula to find the 6th term of the series.

OpenStudy (anonymous):

0.0016?

OpenStudy (anonymous):

an = 625(0.2)^(6 - 1) an = 625(0.2)^(5) an = 625(0.00032) an = 0.2 The 6th term in the series is 0.2.

OpenStudy (anonymous):

How did you get 0.0016? :)

OpenStudy (anonymous):

@TiffaniLynn13 Do you understand how I got 0.2? Or would you like me to explain further?